OpenStudy (anonymous):

rewrite without parentheses and simplify (4u-2)^2

OpenStudy (unklerhaukus):

\[(4u-2)^2=(4u-2)(4u-2)\]\[=4u(4u-2)-2(4u-2)\]\[=16u^2-8u-8u+4\]\[=16u^2-16u+4\]

OpenStudy (anonymous):

^What he said :P